Cost of Units Completed and in Process

The charges to Work in Process—Assembly Department for a period, together with information concerning production, are as follows. All direct materials are placed in process at the beginning of production.

Work in Process—Assembly Department
Bal., 1,600 units, 35% completed 17,440 To Finished Goods, 29,600 units ?
Direct materials, 29,000 units @ $9.50 275,500
Direct labor 84,600
Factory overhead 39,258
Bal. ? units, 45% completed ?

a. Based on the above data, determine the different costs listed below.

1. Cost of beginning work in process inventory completed this period. $
2. Cost of units transferred to finished goods during the period. $
3. Cost of ending work in process inventory. $
4. Cost per unit of the completed beginning work in process inventory, rounded to the nearest cent. $

b. Did the production costs change from the preceding period?

c. Assuming that the direct materials cost per unit did not change from the preceding period, did the conversion costs per equivalent unit increase, decrease, or remain the same for the current period?
Answer #1
Concepts and reason

Cost Accounting: Cost accounting is a method of accounting for recording all the data related to cost. The data is analyzed, summarized and classified in such a way to assist the management in decision making.

Assets: Asset is the resource used by the company to generate income. Assets can be classified into different types, Current assets, long term assets and intangible assets. Assets that are expected to be converted into cash within one year are called as current assets. Assets that are expected to be converted into cash more than one year are called as long-term assets. Asset which cannot be touch and see is called as intangible asset. Goodwill, patents etc. are the examples of intangible assets.

Inventory: Inventory is one of the important current assets of a company. Inventory includes raw materials, work in process and finished goods.

Raw materials: Raw materials inventory is a component used by the company to produce finished goods. Raw material is the basic input used for the production function. Metals, woods, plastics etc. are the examples of raw material inventory.

Conversion cost: Conversion cost is the total of labor and overhead cost of the company.

Fundamentals

Work-in process: It refers to units that could not convert into the finished goods in the current period. Work-in process is also known as semi manufactured products.

Finished goods: Finished goods are the completely finished product ready for sale.

Equivalent units: Partially completed units and fully completed units are expressed in the terms of the fully completed units is called as equivalent units of production.

Beginning work-in process: It refers to the units that are yet to be converted into finished goods. It belongs to the previous period’s closing balance that is shown as beginning balance for the current period.

Ending work-in process: It refers to units that could not convert into the finished goods in the current period.

a.1)

Calculate cost of beginning WIP inventory completed this period.

CostofbeginningWIPinventorycompletedthisperiod}=(Beginningworkinprocess+Conversioncostsincurredduringperiod)=($17,440+(1,040×$4.20))=$17,440+$4,368=$21,808\begin{array}{c}\\\left. \begin{array}{l}\\{\rm{Cost of beginning WIP inventory}}\\\\{\rm{ completed this period}}\\\end{array} \right\}{\rm{ = }}\left( \begin{array}{l}\\{\rm{Beginning work in process }}\\\\{\rm{ + Conversion costs incurred during period}}\\\end{array} \right){\rm{ }}\\\\{\rm{ = }}\left( {{\rm{\$ 17,440 + }}\left( {{\rm{1,040 }} \times {\rm{ \$ 4}}{\rm{.20}}} \right)} \right)\\\\{\rm{ = \$ 17,440 + \$ 4,368 }}\\\\{\rm{ = \$ 21,808}}\\\end{array}

Therefore, the cost of beginning WIP inventory completed this period is $21,808.

Working note:

Calculate cost per equivalent unit.

Particulars
Materials
Completion
1.600
28.000
1,000
100%
100%
Beginning WIP
Units started and completed
Ending WIP
Equivalent

Calculations are given below:

Particulars
Materials
%
Completion
Equivalent
Units
1600
=29600-1600
=1600+29000-29600
Conversion
Completion
1600 0.65
28000

a.2)

Calculate cost of units transferred to finished goods during the period.

Costofunitstransferredtofinishedgoodsduringtheperiod}=((Unitsstartedandcompletedcostsduringtheperiod)+(CostofbeginningWIPinventorycompletedthisperiod))=$383,600+$21,808=$405,408\begin{array}{c}\\\left. \begin{array}{l}\\{\rm{Cost of units transferred to finished}}\\\\{\rm{ goods during the period}}\\\end{array} \right\}{\rm{ = }}\left( \begin{array}{l}\\\left( \begin{array}{l}\\{\rm{Units started and completed }}\\\\{\rm{costs during the period}}\\\end{array} \right){\rm{ }}\\\\{\rm{ + }}\left( \begin{array}{l}\\{\rm{Cost of beginning WIP inventory}}\\\\{\rm{ completed this period}}\\\end{array} \right)\\\end{array} \right){\rm{ }}\\\\{\rm{ = \$ 383,600 + \$ 21,808 }}\\\\{\rm{ = \$ 405,408}}\\\end{array}

Therefore, the cost of units transferred to finished goods during the period is $405,408.

Working note:

Calculate total units started and completed cost during the period is $383,600.

Total
Particulars
Units started and completed (a)
Costs per unit (b)
Units started and completed costs
during current period

a.3)

Calculate cost of ending work in process inventory.

Costofendingworkinprocessinventory}=Materials+Conversion=$9,500+$1,890=$11,390\begin{array}{c}\\\left. \begin{array}{l}\\{\rm{Cost of ending work }}\\\\{\rm{in process inventory}}\\\end{array} \right\}{\rm{ = Materials + Conversion }}\\\\{\rm{ = \$ 9,500 + \$ 1,890 }}\\\\{\rm{ = \$ 11,390}}\\\end{array}

Therefore, the cost of ending work in process inventory is $11,390.

Working note:

Total
Particulars
Ending WIP (a)
Cost per equivalent unit (b)
Cost of ending working process
inventory (a) (b)
Materials
1,00

a.4)

Calculate cost per unit of the completed beginning work in process inventory.

Costperunitofthecompletedbeginningworkinprocessinventory}=CostofbeginningWIPinventoryUnitsofbeginningworkinprocessinventory=$21,8081,600=$13.63\begin{array}{c}\\\left. \begin{array}{l}\\{\rm{Cost per unit of the completed }}\\\\{\rm{beginning work in process inventory}}\\\end{array} \right\}{\rm{ = }}\frac{{{\rm{ Cost of beginning WIP inventory}}}}{{{\rm{Units of beginning work in process inventory}}}}{\rm{ }}\\\\{\rm{ = }}\frac{{{\rm{\$ 21,808}}}}{{{\rm{1,600}}}}{\rm{ }}\\\\{\rm{ = \$ 13}}{\rm{.63}}\\\end{array}

Therefore, the cost per unit of the completed beginning work in process inventory is $13.63.

b)

Yes.

During the period production cost is $13.70.

Working note:

Calculate production cost.

Productioncost=(Costperequivalentunitsofmaterials+Costperequivalentunitsofconversion)=$9.50+$4.20=$13.70\begin{array}{c}\\{\rm{Production cost = }}\left( \begin{array}{l}\\{\rm{Cost per equivalent units of materials }}\\\\{\rm{ + Cost per equivalent units of conversion}}\\\end{array} \right){\rm{ }}\\\\{\rm{ = \$ 9}}{\rm{.50 + \$ 4}}{\rm{.20 }}\\\\{\rm{ = \$ 13}}{\rm{.70}}\\\end{array}

c)

Here, current period conversion cost per equivalent unit is $4.20 and calculated previous period conversion cost is $4.13. Hence, there is an increase in the conversion costs per equivalent unit.

Working note:

Calculate previous period conversion cost.

Previousperiodconversioncost=((Costperunitofthecompletedbeginningworkinprocessinventory)costperequivalentunitsofmaterials)=$13.63$9.50=$4.13\begin{array}{c}\\{\rm{Previous period conversion cost = }}\left( \begin{array}{l}\\\left( \begin{array}{l}\\{\rm{Cost per unit of the completed beginning}}\\\\{\rm{ work in process inventory}}\\\end{array} \right){\rm{ }}\\\\ - {\rm{ cost per equivalent units of materials}}\\\end{array} \right){\rm{ }}\\\\{\rm{ = \$ 13}}{\rm{.63 }} - {\rm{ \$ 9}}{\rm{.50 }}\\\\{\rm{ = \$ 4}}{\rm{.13}}\\\end{array}

Ans: Part a.1

Cost of beginning WIP inventory completed this period is $21,808.

Part a.2

Cost of units transferred to finished goods during the period is $405,408.

Part a.3

Cost of ending work in process inventory is $11,390.

Part a.4

Cost per unit of the completed beginning work in process inventory is $13.63.

Part b

Yes. Production cost is changed.

Part c

Conversion costs per equivalent unit increase for the current period.